OpenStudy (anonymous):

f a ball is thrown straight up into the air with an initial velocity of 40 ft/s, its height in feet after t second is given by y = 40 t - 16 t^2. Find the average velocity for the time period beginning when t = 1 and lasting (i) 0.5 seconds (ii) 0.1 seconds (iii) 0.01 seconds Finally based on the above results, guess what the instantaneous velocity of the ball is when t =1.

OpenStudy (dumbcow):

to find avg velocity, use slope formula \[\frac{f(b) - f(a)}{b-a}\] where a,b are "t" values

OpenStudy (anonymous):

what values do u put in for a,b

OpenStudy (dumbcow):

"b" is ending time .... t = 1.5 or 1.1 or 1.01 "a" is starting time .... t=1

OpenStudy (anonymous):

43.68-24/1.5-1= 39.36 is ths correct??

OpenStudy (dumbcow):

almost how do you get 43.68? did you plug in 1.5 for t?

OpenStudy (anonymous):

yes

OpenStudy (dumbcow):

http://www.wolframalpha.com/input/?i=40t+-16t%5E2%2C+t%3D1.5 should get 24

OpenStudy (anonymous):

sorry my mistake hit plus instead of -

OpenStudy (dumbcow):

so 24-24 = 0 avg velocity for first part is 0

OpenStudy (anonymous):

yes i got that there now

OpenStudy (anonymous):

and what way do i work out the rest ??

OpenStudy (dumbcow):

do same thing, instead of 1.5 use 1.1 and 1.01

OpenStudy (dumbcow):

avg velocity should increase each time and approach 8 which is the instantaneous velocity at t=1

OpenStudy (anonymous):

thanks very much very helpful